OpenStudy (anonymous):

What is the slope of the line between (3, -4) and (-2, 1)? 1 2 -2 -1

OpenStudy (johnweldon1993):

You have 2 points on the line... \[\large (x_1 , y_1) (x_2 , y_2) \] \[\large (3 , -4) (-2 , 1)\] So you plug in those values into the slope formula \[\large m = \frac{y_2 - y_1}{x_2 - x_1}\]

OpenStudy (anonymous):

C?

OpenStudy (johnweldon1993):

Well lets see \[\large m = \frac{y_2 - y_1}{x_2 - x_1} = \frac{1 - (-4)}{-2 - 3} = \frac{5}{-5} = ?\]

OpenStudy (anonymous):

Then its -1?

OpenStudy (johnweldon1993):

Correct :)

OpenStudy (anonymous):

@johnweldon1993 Thank you so much help me with a few more if you dont mind?

OpenStudy (johnweldon1993):

Of course :)

OpenStudy (anonymous):

What type of line is the graph of x = -2? Vertical Horizontal Rising Falling